OpenStudy (anonymous):

A scientist is studying the growth of a particular species of plant. He writes the following equation to show the height of the plant f(n), in cm, after n days: f(n) = 10(1.02)n Part A: When the scientist concluded his study, the height of the plant was approximately 11.04 cm. What is a reasonable domain to plot the growth function? (4 points) Part B: What does the y-intercept of the graph of the function f(n) represent? (2 points) Part C: What is the average rate of change of the function f(n) from n = 1 to n = 5, and what does it represent? (4 points)

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@satellite73 @

OpenStudy (anonymous):

Plz help!

OpenStudy (anonymous):

@ParthKohli

OpenStudy (xhawaiiangirlx):

A. 0 thru 5 (integers) ... (look at part C) B. the starting height of the plant C. f(1) = 10.02 .... f(5) = 11.041 m(ave) = (11.041 - 10.02) / (5 - 1) = 0.255 cm/day this represents the average change of height each day for this time.

OpenStudy (anonymous):

Can I get the explanations please?

OpenStudy (mrnood):

For the firat part you need toknow what reasonable values of n will be plotted on the chart you know that at the end of the trial (i.e after n days) the height is 11.04cm but also that this is 10(1.02)^n so use logs to find n

OpenStudy (anonymous):

ok, im confused

OpenStudy (mrnood):

the trial lasts n days after n days the height is 11.04cm but height = 10(1.02)^n so 11.04 = 10 (1.02)^n use the propoerties of logs and exponents to find n

OpenStudy (anonymous):

oh ok, so n = 5?

OpenStudy (mrnood):

did oyu use the properties of logs and exponents to get that? show yourwork?

OpenStudy (anonymous):

Ok so, 11.04 = 10(1.02)^n 10(1.02)^n = 11.04 10(1.02)^n/10 = 11.04/10 1.02^n = 1.104 log(1.02)^n = log(1.104) n*(log(1.02)) = log(1.104) n = log(1.104)/(log(1.02) n = 4.99 (round to 5)

OpenStudy (anonymous):

ok....

OpenStudy (mrnood):

OK well done now second part the y intercept is when n=0 what does that value mean?

OpenStudy (anonymous):

It represents the starting population. right?

OpenStudy (mrnood):

not the population - since we are measuring height (but I think you've got the right idea) (this exponential function of DOES represent populations! :-)

OpenStudy (anonymous):

Ok so that's part b?

OpenStudy (mrnood):

so wht IS the staring height (it doesnt ask oyu that - but you maybe would lik eot know?)

OpenStudy (anonymous):

Um... .21 cm?

OpenStudy (mrnood):

no - put n= 0 into the equation for height....

OpenStudy (anonymous):

Ok... But the answer cant be true...

OpenStudy (mrnood):

You should also get into th ehabit of doing a 'sanity check' on oyur answers If it is 11cm after 5 days it cannot be 21cm at start. And the answer CAN be true what is 1.02^0??

OpenStudy (anonymous):

1!

OpenStudy (mrnood):

so 10* (1.02)^0 is??

OpenStudy (anonymous):

10

OpenStudy (anonymous):

Im srry, I thought you meant to do the whole equation with 11.04=10(1.02)^0

OpenStudy (mrnood):

good so it starts at 10cm and grows to 11.04 in 5 days the 3rd part says find change of height over 4 days and divide by 4 days so what does that give you?